OpenStudy (anonymous):

if f(x)= 3x^2-4x+1 and g(x)= 3x^2-3 what values are not in the domain of f/g? A. 0 B. 1 C. 3 D. +-1 E. no values to exclude?

OpenStudy (misty1212):

HI

OpenStudy (misty1212):

we need to put \[3x^3-3=0\] and solve we gotta be careful here

OpenStudy (misty1212):

\[3x^2-3=0\\ 3x^2=3\\ x^2=1\\ x=\pm1\]

OpenStudy (anonymous):

Well we have f/g=\[\Large \frac{ 3x^2-4x+1 }{ 3x^2-3 }\]

OpenStudy (misty1212):

say everything but those two numbers

OpenStudy (anonymous):

We also have to look for numbers in the numerator that make it undefined and numbers in the denominator that are undefined or 0(as that will make it like a number divided by 0)

OpenStudy (anonymous):

Are there any numbers that will make the numerator 3x^2-4x+1 undefined?